OpenStudy (df001):

Prove Sec^2ΘCot^2Θ=Csc^2Θ

OpenStudy (df001):

\[\sin ^{2}\theta Cot ^{2}\theta = Csc ^ {2}\theta\]

jimthompson5910 (jim_thompson5910):

is that first term sec^2 ? or sin^2 ?

jimthompson5910 (jim_thompson5910):

I'm going to assume sec^2

jimthompson5910 (jim_thompson5910):

Hint: \[\Large \sec^2(\theta)\cot^2(\theta)=\csc^2(\theta)\] \[\Large \frac{1}{\cos^2(\theta)}*\frac{\cos^2(\theta)}{\sin^2(\theta)}=\csc^2(\theta)\]